MARS DESERT RESEARCH STATION

Journalist Report – May 2nd

Prepared by Victor Roman, Crew Journalist

Date May 2 2017

MDRS Sol 4 update

The sol started very early in the morning. We established a strict schedule of having breakfast, lunch and dinner together so we can get to know each other deeply; so far it’s working perfectly: tosol, we shared thoughts about the current situation in our countries and how corruption harms our economies.

Also each of all us started working in our researches. For example Greenhab Engineer Reyes went to the greenhab and started growing his “uchuvas” (Physalis peruviana); Commander Corrales started recognizing her lab and Crew Engineer Caballero also worked with his rover. By the way, he mentioned a funny anecdote of how security at airports mistakenly took his little robot as a dangerous device.

Later we had an EVA to scout close by places, get used to the suits and vehicles, and take pictures. What we didn’t know is that Health Officer Valencia had a surprise test for us. He made us think he was feeling unwell and we had to evacuate him right away. We were worried for him and we were relieved when we found out it was all set up. Now we are back at the hab having a conversation about our day.

Daily Summary – May 2nd

Crew 180 Daily Summary Report 01052017

MDRS Daily Summary Report for sol (1)

Summary Title (We started the simulation. We had a talk with Shannon. The  laboratory was prepared for  working and  we received the student visitors from a school on earth. Everyone started preparing their project)

Mission Status: (All is fine)

Sol Activity Summary: (We talked to Shannon about the rules on the station. We set the protocols for EVAs and the house. We learned about how to use the EVAs. All the food was counted to be distributed in a better way. We decided to have a shower every 2 or 3 days but just using 8 liters of water. The crew defined Lima-Alpha-Tango-One as the Hab communication password. During our fisrt EVA on SOL-1 we have started at 11:35. We defined the time inside the airlock as one minute before going out)

Look Ahead Plan: (Tomorrow three crew members will have a EVA in the morning. We will start working in the Green Hab and in the laboratory. We will have crew session with Leonardo Valencia  with the aim  of explaining us about their project)

Anomalies in work: (No problems so far)

Weather: Sunny

Crew Physical Status:  The Crew is ok

EVA: (The crew drove along the road till quary Rd in the Special Region Zone

Easting 518500, Northing 4255800 for samples collection of the surface by

the commander Corrales, Y.

The green hab officer Reyes, C. take some pics along the road and helped to the commander to collect the samples.

The EVA Officer Pina, Y. tested the active radiation monitoring and the video-position system through smartphone. The systems are operating adequate.

The crew successfully concluded the EVA at 13:30 and received Earth visitors from Colorado, US to show the hab, lab, and Musk observatory.)

Reports to be file: (Journalist, engineering report and EVA request)

Support Requested: No support

 

Prepared by: Yendry Corrales and Yair Piña

 

Yendry Corrales

Commander